Write a formula for the an in each of the following sequences {an}∞n=1 given by (i) 2,1,4,3,6,5,8,7,... (ii) 1, 3, 6,10,15,... (iii) 1, −4,9,−16,25,−36,... Which ones among these are the subsequence of {n}∞n=1?


1
Expert's answer
2021-01-20T18:07:09-0500

1) an = n - 1, for n=2"i" (even), "i\in\N".

an = n + 1, for n=2"i"-1 (odd), "i\in\N".

2) an = "\frac{n(n+1)}{2}" .

3) an = - n2, for n=2"i" (even), "i\in\N".

an = n2, for n=2"i"-1 (odd), "i\in\N".


{n} = 1, 2, 3, 4, 5...n

1) This sequence can be obtained from {n} without removing members, but with changing the order of the members. She's not a subsequence {n}.

2) This sequence can be obtained from {n} with removing members, without changing the order of the members. She's a subsequence {n}.

3) This sequence cannot be obtained from {n}. She is divergent(dispersent).

Answer: 2.